OpenStudy (anonymous):

how do I solve... cos 2x - cos4x = tan3x tanx(cos2x + cos4x)? I need to prove that is an identity, I do not need to find the value for x

OpenStudy (debbieg):

is it this:\[\cos(2x)-\cos(4x)=\tan(3x)\tan x[\cos(2x)+\cos(4x)]\] are you solving for x?

OpenStudy (anonymous):

yes he is

OpenStudy (anonymous):

I need to prove that is an identity,

OpenStudy (anonymous):

aha

OpenStudy (debbieg):

OK, that's what I thought... "solve" implies something else so threw me off. This looks like a doozy, I'll take a look at it.

OpenStudy (anonymous):

first of all, keep the tan3x tanx terms on right hand side and remaining cos terms to one side, then break the tan3x tanx into (sin3x sinx)/ (cos3x cosx) then apply

OpenStudy (anonymous):

Those identities are off by a factor of two I think...

OpenStudy (anonymous):



OpenStudy (anonymous):

@SandeepReddy The identities should be \[ \sin A\sin B = \frac{1}{2}\left[ \cos(A-b) - \cos(A+B) \right] \]and so on. (Obviously the halves just cancel, but that doesn't make the identities right!) But great trick!

OpenStudy (debbieg):

Slick, @SandeepReddy :)

OpenStudy (debbieg):

Oops, yes good point @Erin001001 ! I see what you mean.

OpenStudy (anonymous):

@Erin001001 you are right at that half part, i almost forgot to include it, thanx for that :)